Hiperspaces homeomorphic to cones Alejandro Illanes, México, México

Keywords: hyperspace, continuum, cone AMS Subject classification: 54B20 Abstract: In this paper we give a step toward the classification of those finite-dimensional metric continua X such that the hyperspace of subcontinua of X is homeomorphic to its topological cone.